Find the charge of a test charge with an electric field of 6.0 x 10⁻⁶ N/C and a force of 0.25 N



1
Expert's answer
2022-02-28T10:06:58-0500
"q=\frac{F}{E}=\rm\frac{0.25\: N}{6.0*10^{-6}\: N/C}=4.2*10^4\: C"